the sum of the square of a positive number and the square of 2 more than the number is 100
Answer by fractalier(6550)

Call the number, x. Then we can write
x^2 + (x+2)^2 = 100
Expand and collect like terms to get
x^2 + x^2 + 4x + 4 = 100
2x^2 + 4x - 96 = 0
divide by two to make things simpler
x^2 + 2x - 48 = 0
now factor
(x + 8)(x - 6) = 0
and
x = -8 and x = 6
but the question asked for a positive number, so
x = 6